Problem 3 (4 point) Determine the x-coordinate of the centroid of the shaded area. 9.14
Answer #1

Xbe b2 x= add aba to both side bx+ay=a?? a%b2-ay? b2 a b-ay? b? Vb-y? x=f(y)=;/62–y? (A= tab;y=0 m = f(y)]’dy a X Area 1 m =

Pls hit a like , if you have any problem let me know in the comment box,?